# Privacy in Speech Technology

### v3

@@ V-A2 Adversarial training
Privacy-preserving neural networks can be trained using adversarial methods, such that side-information in the bottleneck is minimized [48, 64, 92]. Figure 13 illustrates the model structure, where the trusted and threat tasks correspond respectively to utility and privacy, which in turn respectively correspond to the extraction of the target information and side-information. The threat task is independently optimized to extract private side information from the bottleneck output. While keeping the threat task fixed, the encoder and the trusted tasks are jointly optimized to maximize the accuracy [^2] Accuracy refers here to some similarity measure between the message and a corrupt version of the message. of the target information and to minimize the accuracy of private side information.
Adversarial training thus combines maximizing and minimizing the accuracy of private side information. This forces the encoder to minimize private side information passed through the bottleneck since the threat task tries to maximize the extraction of private side information. Similar effects can be achieved by a variety of other formulations, for example, by Siamese training, where the analysis model is specifically optimized to minimize and maximize feature distances between speakers within and across target classes [47]. This approach should cluster extracted information into groups corresponding to the trusted task while minimizing private information.
@@
The principal issue with adversarial training is that theoretical frameworks for certifying performance (e.g. [126]) are not yet widely accepted and there are thus no formal assurances of privacy (cf. section VI-A). The best it can do is to demonstrate the extent to which the chosen adversarial model was unable to extract private side information from the selected category of private information. We can thus expect that some other attacker with a better model could still extract private side-information and that other categories of private information are potentially present in the bottleneck output.
@@
Figure 14: Privacy-preserving speech processing through disentanglement, where speech is decomposed into independent streams of information, each representing a distinct category of information and where the level of anonymization can be individually chosen for each category. Potential points of attack are indicated by red, dashed arrows and exclamation marks.
@@ V-A3 Disentanglement
With disentanglement, we refer to methods that split a speech signal into channels corresponding to distinct categories of information. For example, selecting from table I we can choose to decompose speech into linguistic information, the speaker’s (voice) identity, and emotions [127]. Such disentanglement can be achieved by either or both through adversarial training or by generalizing information bottlenecks to encompass multiple bottleneck channels, where each bottleneck represents a disentangled representation (see fig. 14). Note that provable privacy requires that such channels are statistically independent of each other. That is, each channel represents a distinct and statistically independent attribute of the signal such that the extent of anonymization can be cherry-picked per channel as per use-case e.g. [128, 93, 127, 20].
The privacy-preserving performance of a disentangling system can be readily formulated, provided that channels $k$ are uncorrelated. If the bitrates for private information of respective channels are $B_{k}$, then using the composition theorem, the joint disclosure is $\sum_{k}B_{k}$ [121]. Per-channel disclosures are thus additive such that the design and performance analysis of the privacy-preserving systems is straightforward.
An important challenge with disentanglement is to define the constraints with which information is funneled to the corresponding channel and the methods by which they are made independent of each other. For example, a model can be trained to match specific channels with labels in the dataset (e.g. [93, 127]). Alternatively, using representation learning, we can constrain channels based on objective criteria such as time-scope, i.e. the length of time over which data is integrated (e.g. [129, 118, 130]). Specifically, attributes such as the speaker’s identity are constant over an utterance, while phonemes, intensity and pitch change over a time scale of tens of milliseconds. By modeling the signal in terms of such time scales, we can disentangle attributes that appear on the corresponding time scales.
@@
Figure 15: Protecting privacy by encrypting communication and processing. The gray area indicates the domain that is encrypted and red, double arrows marked with exclamation marks indicate the protected stream of data.
@@ V-B Secure processing
If the transmitted information is encrypted, then it cannot be used for malicious purposes without access to the key. Surprisingly, however, it is possible to process encrypted information when using secure multi-party computation (SMPC), such that the encrypted result of the computation can be returned and opened [131, 132, 133, 108, 134] (see fig. 15). This applies largely to the same attack vectors as Information isolation, including Cloud leak (section IV-D1), False activation (section IV-D2) and Cloud access (section IV-D3), but also WASN authentication (section IV-D4) and Shared device (section IV-D7).
@@
The SMPC umbrella contains several categories of methods such as homomorphic encryption, oblivious transfer, secret sharing, and garbled circuits. The general principle is that in an interaction between two agents, Alice and Bob, Alice can send encrypted data for processing or a request to Bob, such that Bob cannot interpret the input but can perform the requested task. Furthermore, when Alice receives Bob’s response, she cannot deduce the operations that Bob has performed, such that the desired task is completed with mutual privacy preserved. Another category of approaches is distance-preserving hashing techniques, where data is mapped into a space where distances between points are approximately maintained, allowing some operations like clustering, without revealing the actual data [21, 135].
@@
Such secure techniques can be applied, for example, in the extraction of spectral features of speech or speech recognition in the cloud, without plain text access to the speech signal [136, 137]. It can also be applied in privacy-preserving speaker recognition, authentication or verification [31, 39, 41, 43], speech search [138], or for secure sound classification [139]. Finally, we can compare the similarity of audio signals without revealing them [140, 141, 18].
@@
A central problem with such cryptographic protocols is that the floating- or fixed-point operations that can be performed in the encrypted domain are limited, typically to only additions and multiplications [142, 143, 144, 145]. The limitation does not apply to the boolean operations offered by fast fully homomorphic encryption over the torus scheme, garbled circuits, and boolean secret sharing [146]. In contrast, typical speech-processing algorithms also require other floating- or fixed-point operations, such as non-linearities, conditionals, and recursions [147], making the algorithms impossible to implement. While many operations have standard workarounds [148], such as approximation with additions and multiplications, such approximations are often computationally prohibitively complex, see e.g. [136]. In addition, the communication overhead, in terms of the amount of data to send in comparison to the size of the secret message, is very large. For a broader overview of cryptographic methods in speech, see [13].
@@ V-C Privacy-preserving architectures
Perhaps the most obvious solution to protecting privacy, for example in the Cloud leak–scenario, is to limit information flow (to some extent) from the local edge device to the cloud server (see fig. 16). Such limitation or reduction in information flow can be implemented by removing, replacing, or distorting private content as in fig. 16(a) (cf. section V-A). At the extreme, the edge device can be entirely disconnected from the cloud or network, and process information only locally, as in fig. 16(b) [149]. In doing so, we then need to assume that the local device has sufficient capacity and software to complete the requested tasks, that it is not compromised, and that the restriction of information flow is sufficient to ensure that private information is not leaked. Observe, for protecting privacy in scenarios 1 to 4 in fig. 1, local processing is the best protection as nothing is transmitted to a cloud service.
Distributed learning is an approach to training models without the need for centralized data collection [150]. Models are trained on local edge devices and only model updates are shared between the nodes and/or with a central server (see fig. 17). The model updates are, however, calculated from the raw signal and thus describe the original signal in some convoluted way. Though the raw data thus never leaves the edge device in distributed learning, this corresponds to a privacy protection approach where information flow is restricted (cf. section V-A).
Federated learning is one of the flavors of distributed learning, where a central, cloud server collects, merges, and redistributes model updates. It has been used for example in speaker and emotion recognition, language modeling, as well as for unsupervised estimation of microphone clusters in sensor networks [91, 37, 29, 80, 59, 151, 60, 152].
Overall, distributed learning is a promising approach, even if it has several challenges. First, constructing, training, and testing systems architectures is much more complicated than regular machine learning. Second, model updates are model-specific and cannot easily be reused if the model structure is updated. The learning accumulated during training is then effectively lost every time the model is updated, jeopardizing the fair comparison of competing approaches. Third, even if distributed learning does improve privacy, it is not a guarantee for privacy since model updates can also contain private information [153, 150]. To be specific, except for federated learning, distributed learning protects efficiently against threats 1 to 3 in fig. 1, but it may potentially leak private information to other devices, and federated learning may leak information to the cloud.
@@ V-D Acoustic interventions
When a user interacts with a speech interface, then the spoken answer of the speech interface can contain private information. This private information can be overheard by other users in the same acoustic space, and that presents a threat to privacy (see figs. 19 and IV-D5), as in attack vectors False activation (section IV-D2), WASN authentication (section IV-D4), Speech interface leak (section IV-D5), Discussion leak (section IV-D6), and Shared device (section IV-D7). In theory, it would be possible to identify and track users in the same acoustic space and communicate private information only when it does not pose a threat to privacy (see fig. 18). However, tracking itself also poses a threat to privacy, as it is yet another form of private information extracted from the users. This places great trust in the local device’s ability to track and identify the authorization levels of the people present in the acoustic space. Such systems have, however, not yet been widely published.
Figure 19 presents another solution, which uses acoustic interventions to create physical spaces where the private information is, respectively, intelligible or inaudible [154, 155]. This acoustic effect can be achieved either as a sound zone through constructive and destructive interference between loudspeakers (see e.g. [156, 157, 158]) or as an audio spotlight through non-linear interference effects between ultrasound sources [159]. By choosing the spatial location where speech is intelligible and assuming we know where the target user is located, we can thus limit access to private information only to the target speaker.
@@ V-E Improving utility
In many use cases, like False activation (section IV-D2), Speech interface leak (section IV-D4), and Discussion leak (section IV-D6), the actual culprit is the inadequate performance of the service. For example, if a wake word detector is incorrectly triggered (false positive), the system will start to listen to a conversation when it was not supposed to, thus breaching privacy. Since the inadequate wake word detector is thus causing a privacy breach, the best solution is to improve the wake word detector. The treatment thus addresses the cause rather than the symptoms thereof. Preserving privacy thus becomes an issue of systems design and implies using the privacy-by-design approach [160, 161].
This approach has, however, two principal challenges. First, improving utility often requires an increase in computational power and other resource consumption. This is financially costly and has an environmental penalty [162]. In particular, balancing the computational load is easier on a cloud server, and thus, existing resources can be more efficiently used. Second, speech interfaces will always have occasional errors, even with improved utility. That is, the design of privacy-preserving speech technology must include multiple layers of protection, especially when requested operations have large consequences [163, 164]. Say, when the wake word detector is activated, then before any actions which would potentially breach privacy, the system could require that the speaker is in the same room, or the speaker identity is verified, or an extra confirmation step “Are you sure?” or similar. The intrusiveness of such additional protections should then reflect the severity of the potential breach such that the protections are not perceived as overly obtrusive and decrease the utility of the service.
Improving performance can also mean that the system requires additional functionalities. For example, in the case a second local device overhears Alice’s interaction with Bob (Discussion leak) or the primary device (Speech interface leak), the required privacy protection is that the secondary device either 1. is aware that it is not the intended recipient of speech (speech analysis), or 2. notifies the user of its presence (user-interface design).
@@ V-F Privacy-by-design and provable protections
Privacy-by-design is a philosophy or approach to design, where privacy is taken into account from the start of the design process and not as an afterthought [165]. This increases the chances that the system does not violate privacy unnecessarily, since the privacy-preserving performance is considered at every step of the design process.
The privacy-preserving performance of a system can be designed to satisfy design requirements provably, using theoretical frameworks, and differential privacy in particular [166, 167, 13]. Privacy is generally enforced by reducing the accuracy of private information sufficiently, for example, through the addition of noise or by quantization (see fig. 20 and cf. section V-A). To determine how much accuracy must be reduced, differential privacy quantifies how the amount of information available about a speaker’s affiliation to a particular group changes due to an observation. Specifically, suppose an algorithm $M()$ is applied on a set $D$ to give an answer $M(D)=S$. The limit $\epsilon$ for leakage of information (in nits) is then the amount of new information gained through a slightly different dataset $D^{\prime}$ as
@@
It is arguably more informative for users to state the absolute group size within which they remain anonymous. If, through the processing of an observation, an attacker can reduce the group size within which the user remains anonymous to $k$, then we say that the process is $k$-anonymous [168].
Even if individual modules of a system can assure privacy, their interaction in a system can expose privacy. Therefore, the privacy-preservation performance of processes involving speech or other biometric signals must be evaluated on a systems level. Such evaluation uses a concept of biometric reference, which is a data record that stores an attribute of the user, such as the speakers’ identity. New samples are compared to such biometric references to find matching attributes. The three key concepts in such evaluation are [97, 13]
@@
Together, unlinkability and irreversibility thus minimize the amount of information leaked if biometric references are revealed. At the same time, renewability ensures that leaked references can be revoked and new ones created to replace them. Such concepts are central to the legal definitions of privacy and first proposals for their mathematical characterization have emerged [169, 170].
These theoretical design tools and concepts can be used in the design of, say, information bottlenecks (section V-A), to quantify the maximal leak of private information. However, care must be taken to ensure that prior assumptions, such as the equal likelihood of all labels, are met (or enforced) in practical scenarios (cf. [123]). Theoretical methodology for privacy has been applied in, among others, differentially private speech recognition [59], emotion recognition [91], speaker anonymization [53], and federated learning in acoustic sensor network [171]. As theoretical proofs of privacy have obvious advantages, the adoption of this methodology will likely increase.

### v4

@@ V-A2 Adversarial training
Privacy-preserving neural networks can be trained using adversarial methods, such that side-information in the bottleneck is minimized [noe21_interspeech, srivastava19_interspeech, nelus2018itg]. Figure 13 illustrates the model structure, where the trusted and threat tasks correspond respectively to utility and privacy, which in turn respectively correspond to the extraction of the target information and side-information. The threat task is independently optimized to extract private side information from the bottleneck output. While keeping the threat task fixed, the encoder and the trusted tasks are jointly optimized to maximize the accuracy [^2] Accuracy refers here to some similarity measure between the message and a corrupt version of the message. of the target information and to minimize the accuracy of private side information.
Adversarial training thus combines maximizing and minimizing the accuracy of private side information. This forces the encoder to minimize private side information passed through the bottleneck, as the threat task aims to maximize the extraction of such information. Similar effects can be achieved by a variety of other formulations, for example, by Siamese training, where the analysis model is specifically optimized to minimize and maximize feature distances between speakers within and across target classes [nelus19b_interspeech]. This approach should cluster extracted information into groups corresponding to the trusted task while minimizing private information.
@@
The principal issue with adversarial training is that theoretical frameworks for certifying performance (e.g. [NEURIPS2023_e8b0c97b]) are not yet widely accepted and there are thus no formal assurances of privacy (cf. section VI-A). The best it can do is to demonstrate the extent to which the chosen adversarial model was unable to extract private side information from the selected category of private information. We can thus expect that some other attacker with a better model could still extract private side-information and that other categories of private information are potentially present in the bottleneck output.
@@
Figure 14: Privacy-preserving speech processing through disentanglement, where speech is decomposed into independent streams of information, each representing a distinct category of information, and where the level of anonymization can be individually chosen for each category. Potential points of attack are indicated by red, dashed arrows and exclamation marks.
@@ V-A3 Disentanglement
With disentanglement, we refer to methods that split a speech signal into channels corresponding to distinct categories of information. For example, selecting from table I, we can choose to decompose speech into linguistic information, the speaker’s (voice) identity, and emotions [zhou2021vaw]. Such disentanglement can be achieved by either or both through adversarial training or by generalizing information bottlenecks to encompass multiple bottleneck channels, where each bottleneck represents a disentangled representation (see fig. 14). Note that provable privacy requires that such channels be statistically independent of each other. That is, each channel represents a distinct and statistically independent attribute of the signal such that the extent of anonymization can be cherry-picked per channel as per use-case e.g., [aloufi2020privacy, stoidis21_interspeech, zhou2021vaw, williams2022learning].
The privacy-preserving performance of a disentangling system can be readily formulated, provided that channels $k$ are uncorrelated. If the bitrates for private information of respective channels are $B_{k}$, then using the composition theorem, the joint disclosure is $\sum_{k}B_{k}$ [kairouz2015composition]. Per-channel disclosures are thus additive such that the design and performance analysis of the privacy-preserving systems is straightforward.
An important challenge with disentanglement is to define the constraints with which information is funneled to the corresponding channel and the methods by which they are made independent of each other. For example, a model can be trained to match specific channels with labels in the dataset (e.g. [stoidis21_interspeech, zhou2021vaw]). Alternatively, using representation learning, we can constrain channels based on objective criteria such as time-scope, i.e., the length of time over which data is integrated (e.g. [Niekerk2020, neumann2019improving, lotfidereshgi2022cognitive]). Specifically, attributes such as the speaker’s identity are constant over an utterance, while phonemes, intensity, and pitch change over a time scale of tens of milliseconds. By modeling the signal in terms of such time scales, we can disentangle attributes that appear on the corresponding time scales.
@@
Figure 15: Protecting privacy by encrypting communication and processing. The gray area indicates the domain that is encrypted, and the red, double arrows marked with exclamation marks indicate the protected stream of data.
@@ V-B Secure processing
If the transmitted information is encrypted, then it cannot be used for malicious purposes without access to the key. Surprisingly, however, it is possible to process encrypted information when using secure multi-party computation (SMPC), such that the encrypted result of the computation can be returned and opened [yao1986generate, portelo2015logsum, portelo2014privacy, Teixeira2018, cramer2015mpc] (see fig. 15). This applies largely to the same attack vectors as Information isolation, including Cloud leak (section IV-D1), False activation (section IV-D2), and Cloud access (section IV-D3), but also WASN authentication (section IV-D4) and Shared device (section IV-D7).
@@
The SMPC umbrella contains several categories of methods such as homomorphic encryption, oblivious transfer, secret sharing, and garbled circuits. The general principle is that in an interaction between two agents, Alice and Bob, Alice can send encrypted data for processing or a request to Bob, such that Bob cannot interpret the input but can perform the requested task. Furthermore, when Alice receives Bob’s response, she cannot deduce the operations that Bob has performed, such that the desired task is completed with mutual privacy preserved. Another category of approaches is distance-preserving hashing techniques, where data is mapped into a space where distances between points are approximately maintained, allowing some operations like clustering, without revealing the actual data [portelo2015privacy, jimenez2015secure].
@@
Such secure techniques can be applied, for example, in the extraction of spectral features of speech or speech recognition in the cloud, without plain text access to the speech signal [thaine19_interspeech, Zhang2019]. It can also be applied in privacy-preserving speaker recognition, authentication or verification [teixeira22_interspeech, pathak11_interspeech, pathak2012ppsa, portelo2014bprivacy], speech search [portelo2015bprivacy], or for secure sound classification [shashanka2006secure]. Finally, we can compare the similarity of audio signals without revealing them [rane2013privacy, perez2020, zarazaga2022thesis].
@@
A central problem with such cryptographic protocols is that the floating-point or fixed-point operations that can be performed in the encrypted domain are limited, typically to only additions and multiplications [smaragdis2007framework, pathak2012ppsv, pathak2011privacy, pathak2013privacy]. The limitation does not apply to the boolean operations offered by fast fully homomorphic encryption over the torus scheme, garbled circuits, and boolean secret sharing [chillotti2020tfhe]. In contrast, typical speech-processing algorithms also require other floating- or fixed-point operations, such as non-linearities, conditionals, and recursions [chialva2018conditionals], making the algorithms impossible to implement. While many operations have standard workarounds [teixeira2019privacy], such as approximation with additions and multiplications, such approximations are often computationally prohibitively complex, see e.g. [thaine19_interspeech]. In addition, the communication overhead, in terms of the amount of data to send in comparison to the size of the secret message, is very large. For a broader overview of cryptographic methods in speech, see [nautsch2019preserving].
@@ V-C Privacy-preserving architectures
Perhaps the most obvious solution to protecting privacy, for example, in the Cloud leak–scenario, is to limit information flow (to some extent) from the local edge device to the cloud server (see fig. 16). Such limitation or reduction in information flow can be implemented by removing, replacing, or distorting private content as in fig. 16(a) (cf. section V-A). At the extreme, the edge device can be entirely disconnected from the cloud or network, and process information only locally, as in fig. 16(b) [shi2016edge]. In doing so, we then need to assume that the local device has sufficient capacity and software to complete the requested tasks, that it is not compromised, and that the restriction of information flow is sufficient to ensure that private information is not leaked. Observe, for protecting privacy in scenarios 1 to 4 in fig. 1, local processing is the best protection as nothing is transmitted to a cloud service.
Distributed learning is an approach to training models without the need for centralized data collection [liu2021machine]. Models are trained on local edge devices and only model updates are shared between the nodes and/or with a central server (see fig. 17). The model updates are, however, calculated from the raw signal and thus describe the original signal in some convoluted way. Though the raw data thus never leaves the edge device in distributed learning, this corresponds to a privacy protection approach where information flow is restricted (cf. section V-A).
Federated learning is one of the flavors of distributed learning, where a central, cloud server collects, merges, and redistributes model updates. It has been used, for example, in speaker and emotion recognition, language modeling, as well as for unsupervised estimation of microphone clusters in sensor networks [feng22b_interspeech, granqvist20_interspeech, zewoudie2021federated, nelus2021, shoemate2022sottovoce, ro21_interspeech, nguyen2023federated, feng2021attribute].
Overall, distributed learning is a promising approach, even if it has several challenges. First, constructing, training, and testing system architectures is much more complicated than regular machine learning. Second, model updates are model-specific and cannot easily be reused if the model structure is updated. The learning accumulated during training is then effectively lost every time the model is updated, jeopardizing the fair comparison of competing approaches. Third, even if distributed learning does improve privacy, it is not a guarantee for privacy since model updates can also contain private information [Orekondy2018gradientleaks, liu2021machine]. To be specific, except for federated learning, distributed learning protects efficiently against threats 1 to 3 in fig. 1, but it may potentially leak private information to other devices, and federated learning may leak information to the cloud.
@@ V-D Acoustic interventions
When a user interacts with a speech interface, the spoken answer of the speech interface can contain private information. Other users can overhear this private information in the same acoustic space, and that presents a threat to privacy (see figs. 19 and IV-D5), as in attack vectors False activation (section IV-D2), WASN authentication (section IV-D4), Speech interface leak (section IV-D5), Discussion leak (section IV-D6), and Shared device (section IV-D7). In theory, it would be possible to identify and track users in the same acoustic space and communicate private information only when it does not pose a threat to privacy (see fig. 18). However, tracking itself also poses a threat to privacy, as it is yet another form of private information extracted from the users. This places great trust in the local device’s ability to track and identify the authorization levels of the people present in the acoustic space. Such systems have, however, not yet been widely published.
Figure 19 presents another solution, which uses acoustic interventions to create physical spaces where the private information is, respectively, intelligible or inaudible [gardner1998_3daudio, pulkki2018parametric]. This acoustic effect can be achieved either as a sound zone through constructive and destructive interference between loudspeakers (see e.g. [donley2016, nielsen2018, WALLACE2022101285]) or as an audio spotlight through non-linear interference effects between ultrasound sources [yoneyama1983audio]. By choosing the spatial location where speech is intelligible and assuming we know where the target user is located, we can thus limit access to private information only to the target speaker.
@@ V-E Improving utility
In many use cases, like False activation (section IV-D2), Speech interface leak (section IV-D4), and Discussion leak (section IV-D6), the actual culprit is the inadequate performance of the service. For example, if a wake word detector is incorrectly triggered (false positive), the system will start to listen to a conversation when it was not supposed to, thus breaching privacy. Since the inadequate wake word detector is thus causing a privacy breach, the best solution is to improve the wake word detector. The treatment thus addresses the cause rather than the symptoms thereof. Preserving privacy thus becomes an issue of systems design and implies using the privacy-by-design approach [Kalloniatis2009, van2011privacy].
This approach has, however, two principal challenges. First, improving utility often requires an increase in computational power and other resource consumption. This is financially costly and has an environmental penalty [patterson2021carbon]. In particular, balancing the computational load is easier on a cloud server, and thus, existing resources can be more efficiently used. Second, speech interfaces will always have occasional errors, even with improved utility. That is, the design of privacy-preserving speech technology must include multiple layers of protection, especially when requested operations have large consequences [hong2004privacy, backstrom2021edpbcomments]. Say, when the wake word detector is activated, then before any actions which would potentially breach privacy, the system could require that the speaker is in the same room, or the speaker’s identity is verified, or an extra confirmation step “Are you sure?” or similar. The intrusiveness of such additional protections should then reflect the severity of the potential breach, such that the protections are not perceived as overly obtrusive and decrease the utility of the service.
Improving performance can also mean that the system requires additional functionalities. For example, in the case of a second local device overhearing Alice’s interaction with Bob (Discussion leak) or the primary device (Speech interface leak), the required privacy protection is that the secondary device either 1. is aware that it is not the intended recipient of speech (speech analysis), or 2. notifies the user of its presence (user-interface design).
@@ V-F Privacy-by-design and provable protections
Privacy-by-design is a philosophy or approach to design, where privacy is taken into account from the start of the design process and not as an afterthought [schaar2010privacy]. This increases the chances that the system does not violate privacy unnecessarily, since the privacy-preserving performance is considered at every step of the design process.
The privacy-preserving performance of a system can be designed to satisfy design requirements provably, using theoretical frameworks, and differential privacy in particular [dwork2006differential, dwork2008differential, nautsch2019preserving]. Privacy is generally enforced by reducing the accuracy of private information sufficiently, for example, through the addition of noise or by quantization (see fig. 20 and cf. section V-A). To determine how much accuracy must be reduced, differential privacy quantifies how the amount of information available about a speaker’s affiliation to a particular group changes due to an observation. Specifically, suppose an algorithm $M()$ is applied on a set $D$ to answer $M(D)=S$. The limit $\epsilon$ for leakage of information (in nits) is then the amount of new information gained through a slightly different dataset $D^{\prime}$ as
@@
It is arguably more informative for users to state the absolute group size within which they remain anonymous. If, through the processing of an observation, an attacker can reduce the group size within which the user remains anonymous to $k$, then we say that the process is $k$-anonymous [sweeney2002k].
Even if individual modules of a system can assure privacy, their interaction in a system can expose privacy. Therefore, the privacy-preservation performance of processes involving speech or other biometric signals must be evaluated on a systems level. Such evaluation uses a concept of biometric reference, which is a data record that stores an attribute of the user, such as the speaker’s identity. New samples are compared to such biometric references to find matching attributes. The three key concepts in such evaluation are [ISOIEC224745_2022, nautsch2019preserving]
@@
Together, unlinkability and irreversibility thus minimize the amount of information leaked if biometric references are revealed. At the same time, renewability ensures that leaked references can be revoked and new ones created to replace them. Such concepts are central to the legal definitions of privacy, and first proposals for their mathematical characterization have emerged [vauquier25_interspeech, cohen2020towards].
These theoretical design tools and concepts can be used in the design of, say, information bottlenecks (section V-A), to quantify the maximal leak of private information. However, care must be taken to ensure that prior assumptions, such as the equal likelihood of all labels, are met (or enforced) in practical scenarios (cf. [vali2024porcupine]). Theoretical methodology for privacy has been applied in, among others, differentially private speech recognition [shoemate2022sottovoce], emotion recognition [feng22b_interspeech], speaker anonymization [shamsabadi2023diff], and federated learning in an acoustic sensor network [el2022differential]. As theoretical proofs of privacy have obvious advantages, the adoption of this methodology will likely increase.